help me find the derivation of Vm=
(i want to find the Vm)
Soave-Redlich-Kwong (SRK) EOS RT aa Vm - b Vm (Vm + b)
Answer #1

RT PE Vm-b αα Vm(Vm + b)

P RT Vm+ b Vm - 6 Vm +b aa Vm - b Vm (Vm +b) Vm - 6

RT(Vm +b) 72 - 12 >> P = aa (Vm - b) V (V2 - 62)

p= VmRT(Vm + b) - aa (Vm - b) V2 - 62

= P(V2 - 12) = VmRT (Vm+b) - aa (Vm - b)

=pV2-pb2 = RTV+bRTVm - aaVm + aab

=vp - RT) + Vm(aa - bRT) + (-pb2 - aab) = 0

Which is quadratic in V_m, Hence

-aa - bRT) + (aa - bRT)2 - 4(p - RT)-pb2 - aab)) 2p - RT)